OpenStudy (anonymous):

what is the number that is one half of one quarter of one tenth of 400?

zepdrix (zepdrix):

This isn't too bad if you do it piece by piece. one tenth of 400 is?

OpenStudy (jdoe0001):

\(\bf half\quad \cfrac{1}{2}\quad quarter\quad \cfrac{1}{4}\quad tenth\cfrac{1}{10}\quad and\quad 400 \\ \quad \\ \textit{half of a quarter}\to \cfrac{1}{4}\div \cfrac{1}{2}\to \cfrac{1}{4}\cdot \cfrac{2}{1} \\ \quad \\ \textit{one tenth of that is}\to \cfrac{1}{4}\div \cfrac{1}{2}\div \cfrac{1}{10}\to \cfrac{1}{4}\cdot \cfrac{2}{1}\cdot \cfrac{10}{1} \\ \quad \\ \textit{of 400 will be }\cfrac{1}{4}\cdot \cfrac{2}{1}\cdot \cfrac{10}{1}\cdot 400\)

OpenStudy (jdoe0001):

actually.. shoot.. lemme fix that some =)

zepdrix (zepdrix):

ya D:

OpenStudy (jdoe0001):

heheh

zepdrix (zepdrix):

it was so pretty though lol

OpenStudy (jdoe0001):

\(\bf half\quad \cfrac{1}{2}\quad quarter\quad \cfrac{1}{4}\quad tenth\cfrac{1}{10}\quad and\quad 400 \\ \quad \\ \textit{half of a quarter}\to \cfrac{1}{4}\cdot \cfrac{1}{2} \\ \quad \\ \textit{one tenth of that is}\to \cfrac{1}{4}\cdot \cfrac{1}{2}\cdot \cfrac{1}{10} \\ \quad \\ \textit{of 400 will be }\cfrac{1}{4}\cdot \cfrac{1}{2}\cdot \cfrac{1}{10}\cdot 400\)

OpenStudy (anonymous):

Tanx

OpenStudy (jdoe0001):

yw